# Kinetically Consistent Coarse Graining using Kernel-based Extended Dynamic Mode Decomposition

### v1

@@ Abstract
In this paper, we show how kernel-based approximation to the Koopman generator – the kgEDMD algorithm – can be used to identify implied timescales and meta stable sets in stochastic dynamical systems, and to learn a coarse-grained dynamics on reduced variables, which retains the essential kinetic properties of the full model. The centerpiece of this study is a learning method to identify an effective diffusion in coarse-grained space by leveraging the kgEMD model for the Koopman generator. By combining this method with force matching, a complete model for the effective dynamics can be inferred. Using a two-dimensional model system and molecular dynamics simulation data of alanine dipeptide, we demonstrate that the proposed method successfully and robustly recovers the essential theormodynamic and kinetic properties of the full model. The parameters of the method can be determined using standard model validation techniques.
@@ 1 Introduction
Stochastic simulations of large-scale dynamical systems are widely used to model the behaviour of complex systems, with applications in computational physics, chemistry, materials science, and engineering. Many examples of such systems are high dimensional and subject to meta-stability, which means the system remains trapped in a set of geometrically similar configurations, while transitions to another such state are extremely rare. As a consequence, it becomes necessary to produce very long simulations in order to make statistically robust predictions. A prime example are atomistic molecular dynamics simulations (MD) 1 of macro-molecules, where meta-stability is typically caused by high energetic barriers separating deep potential energy minima 2 . As a result, it requires specialized high-performance computing facilities to reach the required simulation times, or it may just not be feasible at all 3, 4 .
@@
Coarse graining (CG) describes the process of replacing the original dynamical system by a surrogate model on a (much) lower-dimensional space of descriptors 5, 6 , in such a way that certain properties of the original dynamics are preserved by the surrogate model. CG models can enable scientists to achieve much longer simulation times because of the reduced computational cost, while maintaining predictive capabilities of the full-order model. Setting up a CG model typically requires the following steps: first, the choice of a linear or non-linear mapping (CG map) from full state space to a lower-dimensional space, where the latter serves as the state space of the surrogate model. Second, definition of a parametric model class for the surrogate dynamics. Finally, determination of the parameters for that model class.
@@
The first step is crucial to the CG model’s success, and has been a very active area of research for a long time, see Refs. 7, 8, 4 for reviews on this topic. In this study, we are concerned with the second and mainly the third step. CG models have often been parameterized using physically intuitive functional forms of, for instance, the coarse-grained energy. More recently, much more general functional forms have been used for the CG parameters, which are then approximated by powerful model classes, such as deep neural networks or reproducing kernels 9, 10, 11 , which is the approach we follow in this paper. We focus on CG for reversible stochastic differential equations (SDE) with a Boltzmann-type invariant distribution, which includes most popular simulation engines in MD, e.g. Langevin dynamics. Following the projection approach from Refs. 12, 13 , we likewise parameterize the coarse-grained model as a reversible SDE, disregarding memory terms.
@@
Concerning parameter selection, the success of machine learning (ML) in recent years has led to the development of many powerful learning schemes for the parameters of a CG model. Examples are free energy learning 14 , and force matching 15 , among others. Many of these learning methods are geared towards ensuring thermodynamic consistency, which means that the surrogate model is trained to sample the marginalized Boltzmann distribution in CG space, thus ensuring accurate estimation of average quantities. Ensuring faithful reproduction of kinetic properties, such as time-correlation functions or transition time scales, is a much less developed topic, see e.g. 16, 17 for prior works in this area. In this paper, we focus on recovery of the transition time scales associated to meta stable states.
@@
Transition rates and time scales are directly related to the leading spectrum of the system’s transfer or Koopman operator 18, 19, 20, 21 , which is the linear operator propagating expectation values of observable functions in time. Equivalently, one may also consider the spectrum of the associated Koopman generator (Kolmogorov operator for SDEs) close to zero. This connection has been at the heart of the Markov state modeling (MSM) approach 22, 23, 24 and many important developments based on it 25, 26, 27 . The spectral matching approach 28 makes use of this connection, by first parameterizing the CG model as a linear expansion using a fixed set of basis functions, and then solving a regression problem to recover the eigenvalues of the Koopman generator. This idea was formalized in Ref. 29 , by suggesting to regress on a full matrix representation of the Koopman generator, to capture as much kinetic content as possible. The generator matrix can be estimated a priori by a data-driven algorithm called generator EDMD (gEDMD).
In this study, we significantly improve the generator-based spectral matching, by formulating the approach using kernel methods. Reproducing kernels 30, 31 offer powerful model classes which have been applied in many areas of machine learning. Kernel-based approximation of the Koopman generator (kgEDMD) was first described in Ref. 32 We use the kgEDMD matrix and a (reduced and whitened) basis of kernel functions to formulate a learning problem for a generally state-dependent effective diffusion, which optimally approximates the Koopman generator. Our approach only requires simulation data of the full system, the choice of the kernel function, and measurements of the local diffusion, which is analogue to the local mean force in force matching. Combined with the latter, this approach provides complete access to the effective dynamics associated to a reversible system.
The key contributions of this work are:
  • We formulate a data-based learning problem for the effective diffusion of a coarse-grained SDE. The only hyper-parameters to be tuned are those of the kernel function. The method is robust to statistical noise and ill-conditioning as it is based on a whitened and truncated kernel bases.
  • We show that diffusion learning and force matching can be combined into a single pipeline to completely parameterize a coarse-grained SDE for reversible systems.
@@
  • We show that thermodynamic and kinetic consistency are achieved by the method using two test cases, a two-dimensional model system and molecular dynamics simulations of the alanine dipeptide.
@@
The structure of the paper is as follows: we introduce the required background on SDEs, coarse graining, and Koopman operator learning in Section 2. Our kernel-based learning framework is then presented in Section 3, while the numerical examples follow in Section 4. An overview of the notation is provided in Table 1, supplementary information on simulation details and model selection is given in the Appendix.
@@ 2.1 Stochastic processes
$a(x)=\sigma^{T}(x)\sigma(x).$ (2)
Processes like Equation (1) are referred to as diffusion processes. A standard example commonly used in molecular modeling is overdamped Langevin dynamics
@@
where $F:\Omega\rightarrow\mathbb{R}$ is the potential energy, $\beta=(k_{B}T)^{-1}$ and $\gamma$ are constants corresponding to the inverse temperature and the friction, respectively. The invariant measure for $X_{t}$ in Equation (3) is the Boltzmann distribution $\mu\propto\exp(-\beta F)$, and the dynamics are reversible with respect to $\mu$. More generally, a reversible SDE with invariant measure $\mu\propto\exp(-F)$ can be parameterized in terms of the generalized scalar potential $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, and the diffusion covariance $a$, as follows 33
@@
We will mostly consider reversible SDEs in this paper, and make use of the parameterization in Equation (4) when formulating learning methods.
2.2 Koopman and its generator
Koopman theory 34, 35 lifts the dynamics in Equation (1) into an infinite-dimensional space of observable functions to express the dynamics linearly. More precisely, the family of Koopman operators ${\mathcal{K}^{t}}$ for stochastic dynamics is defined as
@@ 2.2 Koopman and its generator
where $\mathbb{E}\left[\cdot\right]$ denotes the expected value. The operators $\mathcal{K}^{t}$ form a semigroup, that is $\mathcal{K}^{s+t}=\mathcal{K}^{s}\mathcal{K}^{t}$, and one can define their infinitesimal generator $\mathcal{L}$ as the time-derivative of the expectation value
$\mathcal{L}\psi(x)=\frac{d}{dt}\mathbb{E}^{x}\left[\psi(x_{t})\right]|_{t=0}.$ (6)
Stochastic calculus shows that for smooth observable functions, the generator can be written as 33
@@
where $a$ and $b$ are the diffusion and drift terms defined above, $\nabla^{2}[\cdot]$ is the Hessian matrix of a function, and the colon :::: is a short-hand for the dot product between two matrices, also called Frobenius inner product. For overdamped Langevin dynamics, Eq. (7) simplifies to
@@
2.3 Spectral decomposition
A key quantity of interest are the eigenvalues and eigenfunctions of the generator. The study of spectral components of the generator helps us identify the long-time dynamics of the system. In molecular dynamics, we expect to find a number of eigenvalues close to zero, followed by a spectral gap. These low-lying eigenvalues are indicating the number of meta stable states of the system, which are the macrostates the system stays in the longest 18 . We write the eigenvalue problem for the generator as
@@ 2.3 Spectral decomposition
The eigenvalues $\lambda_{i}$ of $-\mathcal{L}$ must be non-negative, and the lowest eigenvalue $\lambda_{1}=0$ is non-degenerate under very mild conditions 36 $0=\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq...$ We also refer to the eigenvalues as rates, and to their reciprocals as implied timescales 22
@@
2.4 Coarse graining
One of the main motivations of this work is to learn an SDE representing the full dynamics (1) on a coarse-grained space. Coarse graining (CG) is realized by mapping the state space $\Omega$ onto a lower-dimensional space $\hat{\Omega}\subset\mathbb{R}^{d_{\xi}}$ by means of a smooth CG function $\xi$. We write $\nu$ for the marginal distribution of the full-space invariant measure $\mu$ projected onto CG space. An important object is the space $L^{2}_{\nu}$ of functions on $\hat{\Omega}$ which are square integrable with respect to the weight function $\nu$. The space $L^{2}_{\nu}$ is a subspace of $L^{2}_{\mu}$ comprising all functions that depend only on the value of $\xi$. Moreover, the orthogonal projector onto $L^{2}_{\nu}$ from $L^{2}_{\mu}$ is the conditional expectation operator 12, 13
@@ 2.4 Coarse graining
where $z$ is a position in CG space. This operator calculates the average of a function $\psi$ over all $x\in\Omega$ whose projection onto CG space is the same point $z\in\hat{\Omega}$.
Considering $\mathcal{P}$, following the exposition in 13, one can define the projected generator
@@
It turns out its action on a function $\varphi\in L^{2}_{\nu}$ is given by
@@
and the practical task of coarse graining is to approximate them numerically. An important result, shown in 13 , and that we will use later on, is that the inner product of the projected generator with respect to the invariant measure $\nu$ in CG space is equal to the inner product of the original generator with respect to the invariant measure $\mu$ of the full space
$\left\langle\mathcal{L}\psi\circ\xi,\,\varphi\circ\xi\right\rangle_{\mu}=\left\langle\mathcal{L}^{\xi}\psi,\,\varphi\right\rangle_{\nu}$ (15)
where $\psi$ and $\varphi$ are functions in $L^{2}_{\nu}$.
2.5 Dynamic mode decomposition and its extension
Dynamic Mode Decomposition (DMD) is a conceptually simple method that provides a linear map to propagate the state of a dynamical system in time. The method was originally proposed in Ref. 37 and since then, many variations and extensions have been proposed, see Ref. 38, 39 for an overview. Basically, DMD looks for the best-fit linear operator $\mathbf{A}$ that advances the system state by solving the minimization problem in Equation (16)
$\mathbf{A}=\operatorname*{arg\,min}_{\mathbf{A}^{\prime}}\|\mathbf{X}^{\tau}-\mathbf{A}\mathbf{X}\|_{F}=\mathbf{X}^{\tau}\mathbf{X}^{\dagger}$ (16)
where $\mathbf{X}$ and $\mathbf{X}^{\tau}$ are collections of time-shifted snapshots of the dynamics, and $\dagger$ denotes the pseudo-inverse. Although DMD is a popular and widely-used method in data-driven modeling, it can only provide a linear model and the systems we are interested in for molecular dynamics are highly nonlinear.
Improved results can be obtained by using nonlinear functions instead of just the raw state variables. This method is known as Extended DMD (EDMD) 40 , that can provide more accurate models by capturing nonlinearities. After defining a dictionary of observables, $\{\psi_{i}\}_{i=1}^{n}$ (that are functions of the state space $\Omega$), we compile the evaluations of all dictionary functions at all snapshots into matrices
$\mathbf{\Psi}$ $=\left[\psi_{i}(x_{l})\right]_{i,l}\in\mathbb{R}^{n\times m},$ $\mathbf{\Psi}^{\tau}$ $=\left[\psi_{i}(x_{l+1})\right]_{i,l}\in\mathbb{R}^{n\times m},$
where we assume for simplicity that subsequent snapshots are separated by time $\tau$. We then solve the optimization problem:
$\mathbf{K}=\operatorname*{arg\,min}_{\mathbf{K^{\prime}}}\|\mathbf{\Psi}^{\tau}-\mathbf{K}^{\prime}\mathbf{\Psi}\|_{F}=\mathbf{\Psi}^{\tau}\mathbf{\Psi}^{\dagger}.$ (17)
The solution of Equation (17) can be obtained numerically using singular value decomposition. In the limit of infinite data, EDMD constitutes a finite-dimensional projection (Galerkin projection) of the Koopman operator onto the subspace spanned by the dictionary functions 40 .
@@ 2.6 Generator EDMD
The infinitesimal generator of the Koopman operator can also be approximated by an algorithm similar to EDMD, called gEDMD 29 . Once again, given a set of scalar basis functions $\psi(x)=\{\psi_{1}(x),...,\psi_{n}(x)\}$, and training data $\{x_{l}\}_{l=1}^{m}$, we form the matrices
@@
using the analytical formula (7) to evaluate the second of these matrices. Solving an analogous regression problem to (17) leads to the matrix
$\mathbf{L}^{\top}=(\mathcal{L}\mathbf{\Psi})\mathbf{\Psi}^{\dagger}=\hat{\mathbf{A}}\hat{\mathbf{G}}^{-1},$ (18)
@@
$\begin{gathered}\hat{\mathbf{A}}_{ij}=\frac{1}{m}\sum_{k=1}^{m}(\mathcal{L}\psi_{i})(x_{k})\psi_{j}(x_{k}),\\ \hat{\mathbf{G}}_{ij}=\frac{1}{m}\sum_{k=1}^{m}\psi_{i}(x_{k})\psi_{j}(x_{k}).\end{gathered}$ (19)
As for EDMD, it was proved in Ref. 29 that in the infinite-data limit, gEDMD converges to the Galerkin projection of the generator onto the space spanned by the basis functions
$\mathbf{L}^{\top}$ $=\mathbf{A}\mathbf{G}^{-1},$ $\mathbf{A}_{ij}$ $=\left\langle\mathcal{L}\psi_{i},\,\psi_{j}\right\rangle_{\mu}$ $\mathbf{G}_{ij}$ $=\left\langle\psi_{i},\,\psi_{j}\right\rangle_{\mu}.$ (20)
The matrices $\mathbf{G}$ and $\mathbf{A}$ are often called mass and stiffness matrix, respectively. For arbitrary stochastic dynamics, the computation of $\mathbf{A}$ involves a second-order differentiation as shown in Equation (7). However, if the stochastic dynamics are reversible, only first-order derivatives are required to compute the matrix $\mathbf{A}$, as the generator satisfies the following integration-by-parts formula
$\mathbf{A}_{ij}=\left<\mathcal{L}\psi_{i},\psi_{j}\right>_{\mu}=-\frac{1}{2}\int\nabla\psi_{i}\sigma\sigma^{T}\nabla\psi^{T}_{j}\,\mathrm{d}\mu.$ (21)
If the basis functions are actually defined in a CG space $\hat{\Omega}$, that is $\psi_{i}(x)=\psi_{i}(\xi(x))$, then by the chain rule the matrix $A$ can be written as
$\begin{split}\mathbf{A}_{ij}=-\frac{1}{2}\int\nabla_{x}\psi_{i}\sigma\sigma^{T}\nabla_{x}\psi^{T}_{j}\,\mathrm{d}\mu=-\frac{1}{2}\int(\nabla_{z}\psi_{i}\nabla_{x}\xi)\sigma\sigma^{T}(\nabla_{x}\xi^{T}\nabla_{z}\psi^{T}_{j})\,\mathrm{d}\mu\\ =-\frac{1}{2}\int(\nabla_{z}\psi_{i})(\nabla_{x}\xi\sigma\sigma^{T}\nabla_{x}\xi^{T})(\nabla_{z}\psi^{T}_{j})\,\mathrm{d}\mu.\end{split}$ (22)
@@
$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi\sigma\sigma^{T}\nabla_{x}\xi^{T}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ (23)

### v2

@@ Abstract
In this paper, we show how kernel-based models for the Koopman generator – the gEDMD method – can be used to identify coarse-grained dynamics on reduced variables, which retain the slowest transition timescales of the original dynamics. The centerpiece of this study is a learning method to identify an effective diffusion in coarse-grained space, which is similar in spirit to the force matching method. By leveraging the gEDMD model for the Koopman generator, the kinetic accuracy of the CG model can be evaluated. By combining this method with a suitable learning method for the effective free energy, such as force matching, a complete model for the effective dynamics can be inferred. Using a two-dimensional model system and molecular dynamics simulation data of alanine dipeptide and the Chignolin mini-protein, we demonstrate that the proposed method successfully and robustly recovers the essential kinetic and also thermodynamic properties of the full model. The parameters of the method can be determined using standard model validation techniques.
@@ 1 Introduction
Stochastic simulations of large-scale dynamical systems are widely used to model the behaviour of complex systems, with applications in computational physics, chemistry, materials science, and engineering. Many examples of such systems are high dimensional and subject to meta-stability, which means the system remains trapped in a set of geometrically similar configurations, while transitions to another such state are extremely rare. As a consequence, it becomes necessary to produce very long simulations in order to make statistically robust predictions. A prime example are atomistic molecular dynamics simulations (MD) 1 of macro-molecules, where meta-stability is typically caused by high energetic barriers separating deep potential energy minima 2 . As a result, it requires specialized high-performance computing facilities to reach the required simulation times, or it may just not be feasible at all 3 .
@@
Coarse graining (CG) describes the process of replacing the original dynamical system by a surrogate model on a (much) lower-dimensional space of descriptors 4, 5 , in such a way that certain properties of the original dynamics are preserved by the surrogate model. CG models can enable scientists to achieve much longer simulation times because of the reduced computational cost, while maintaining predictive capabilities of the full-order model. Setting up a CG model typically requires the following steps: first, the choice of a linear or non-linear mapping (CG map) from full state space to a lower-dimensional space, where the latter serves as the state space of the surrogate model. Second, definition of a parametric model class for the surrogate dynamics. Finally, determination of the parameters for that model class.
@@
The first step is crucial to the CG model’s success, and has been a very active area of research for a long time, see Refs. 6, 7, 8 for reviews on this topic. In this study, we only show examples of low-dimensional CG coordinates that have already been validated, while the problem of learning high-dimensional and fully transferrable CG models is left for future studies. Our focus is instead on the second and mainly the third step. CG models have often been parameterized using physically intuitive functional forms for the coarse-grained energy. More recently, much more general functional forms have been used for the CG parameters, which are then approximated by powerful model classes, such as deep neural networks or reproducing kernels 9, 10, 11 , which is the approach we follow in this paper. We study CG for reversible stochastic differential equations (SDE) with a Boltzmann-type invariant distribution, which is applicable to most popular simulation engines in MD, e.g. Langevin dynamics. Following the projection approach from Refs. 12, 13 , we likewise parameterize the coarse-grained model as a reversible SDE, disregarding memory terms.
@@
The success of machine learning (ML) in recent years has led to the development of many powerful learning schemes for the parameters of a CG model. Examples are free energy learning 14 , and force matching 15 , among others. Many of these learning methods are geared towards ensuring thermodynamic consistency, which means that the surrogate model is trained to sample the marginalized Boltzmann distribution in CG space, thus ensuring accurate estimation of average quantities. Ensuring faithful reproduction of kinetic properties, such as time-correlation functions or transition timescales, is a much less developed topic, see Ref. 16 for an overview of prior work in this area. In this paper, we focus on the recovery of implied transition timescales 17 associated to meta stable states.
@@
Implied transition timescales are defined in terms of the leading spectrum of the system’s transfer or Koopman operator 18, 19, 20, 21 . Equivalently, one may also consider the spectrum of the associated Koopman generator (Kolmogorov operator for SDEs) close to zero. This connection has been at the heart of the Markov state modeling (MSM) approach 17, 22, 23 and many important developments based on it 24, 25, 26 . The spectral matching approach 27 was the first to make use of this connection, by first parameterizing the CG model as a linear expansion of fixed basis functions, and then solving a regression problem to recover the eigenvalues of the Koopman generator. This idea was formalized in Ref. 28 , by suggesting to regress on a full matrix representation of the Koopman generator. The generator matrix can be estimated a priori by a data-driven algorithm called generator EDMD (gEDMD).
In this study, we significantly improve on the idea of leveraging the Koopman generator for the identification of coarse grained models. Our key contributions are the following:
  • We formulate a data-based learning problem for the effective diffusion of a coarse-grained SDE. This formulation is analogous to the force matching approach for the coarse-grained energy. Just as force matching relies on measurements of the local mean force, our approach rests on a similar quantity called local diffusion.
  • We suggest to parametrize the diffusion by a basis of random Fourier features 29 , which form a widely-used approximation technique for reproducing kernels. Random features offer a compromise between representational power and computational efficiency. The only hyper-parameters to be tuned are those of the kernel function. The method is robust to statistical noise and ill-conditioning as it is based on a whitened and truncated basis set.
  • We show how the gEDMD method can be leveraged to evaluate the kinetic consistency of the learned CG model by comparing its eigenvalues to those of the reference gEDMD matrix. This assessment does not require simulations of the CG model.
@@
  • We show that kinetic and also theromodynamic consistency are achieved by the method using three test cases, a two-dimensional model system and molecular dynamics simulations of the alanine dipeptide and the Chignolin mini-protein. For the molecular systems, we consider CG models in low-dimensional reaction coordinate spaces, and we employ an overdamped assumption to simplify the learning process.
@@
The structure of the paper is as follows: we introduce the required background on SDEs, coarse graining, and Koopman operator learning in Section 2. Our learning framework is then presented in Section 3, while the numerical examples follow in Section 4. An overview of the notation is provided in Table 1, supplementary information on simulation details and model selection is given in the Appendix.
@@ 2.1 Stochastic processes
$a(x)=\sigma(x)\sigma^{\top}(x).$ (2)
A standard example for Eq. (1), commonly used in molecular modeling, is overdamped Langevin dynamics
@@
where $F:\Omega\rightarrow\mathbb{R}$ is the potential energy, $\beta=(k_{B}T)^{-1}$ and $\gamma$ are constants corresponding to the inverse temperature and the friction, respectively. The invariant measure for $X_{t}$ in Equation (3) is the Boltzmann distribution $\mu\propto\exp(-\beta F)$, and the dynamics are reversible with respect to $\mu$. More generally, a reversible SDE with invariant measure $\mu\propto\exp(-F)$ can be parameterized in terms of the generalized scalar potential $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, and the diffusion covariance $a$, as follows 30
@@
We will only consider reversible SDEs in this paper, and make use of the parametrization in Equation (4) when formulating learning methods.
2.2 Koopman generator and spectral decomppsition
Koopman theory 31, 32 lifts the dynamics in Equation (1) into an infinite-dimensional space of observable functions to express the dynamics linearly. More precisely, the family of Koopman operators ${\mathcal{K}^{t}}$ for stochastic dynamics is defined as
@@ 2.2 Koopman generator and spectral decomppsition
where $\mathbb{E}\left[\cdot\right]$ denotes the expected value. The associated infinitesimal generator $\mathcal{L}$ is the time-derivative of the expectation value, which can be written as a linear differential operator:
@@
where $a$ and $b$ are the diffusion and drift terms defined above, $\nabla^{2}[\cdot]$ is the Hessian matrix of a function, and the colon :::: is a short-hand for the dot product between two matrices. For overdamped Langevin dynamics, Eq. (6) simplifies to
@@
The key quantity of interest are the eigenvalues and eigenfunctions of the generator. The study of spectral components of the generator helps us identify the long-time dynamics of the system. In molecular dynamics, we expect to find a number of eigenvalues close to zero, followed by a spectral gap. These low-lying eigenvalues are indicating the number of meta-stable states of the system, which are the macro states the system stays in the longest 18 . We write the eigenvalue problem for the generator as
@@
The eigenvalues $\lambda_{i}$ of $-\mathcal{L}$ must be non-negative, and the lowest eigenvalue $\lambda_{1}=0$ is non-degenerate 33 : $0=\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq...$ We also refer to the eigenvalues as rates, and to their reciprocals as implied timescales 17
@@
2.3 Coarse graining and projection
One of the main motivations of this work is to learn an SDE representing the full dynamics (1) on a coarse grained space. Coarse graining (CG) is realized by mapping the state space $\Omega$ onto a lower-dimensional space $\hat{\Omega}\subset\mathbb{R}^{d}$ by means of a smooth CG function $\xi$. We write $\nu~{}\propto\exp(-\beta F^{\xi})$ for the marginal distribution of the full-space invariant measure $\mu$, where $F^{\xi}$ is the free energy in the CG space.
To define dynamics in the CG space, we use the conditional expectation operator 12, 13
@@ 2.3 Coarse graining and projection
where $z$ is a position in CG space. This operator calculates the average of a function $\psi$ over all $x\in\Omega$ whose projection onto CG space is the same point $z\in\hat{\Omega}$. Following the exposition in 13 , one can define the projected generator
@@
It turns out its action on a function $\varphi=\varphi(z)$ in CG space is given by
@@
and the practical task of coarse graining is to approximate them numerically.
@@ 2.4 Generator EDMD
Numerical approximations to the infinitesimal generator $\mathcal{L}$ can be obtained by a data-driven learning method called generator extended dynamic mode decomposition  28(gEDMD). Given a finite set of scalar basis functions $\psi(x)=\{\psi_{1}(x),...,\psi_{n}(x)\}$, and training data $\{x_{l}\}_{l=1}^{m}$ sampled from the invariant measure $\mu$, we form the matrices
@@
using the analytical formula (6) to evaluate the second of these matrices. The solution of a linear regression problem leads to the matrix approximation
$\mathbf{L}=\hat{\mathbf{G}}^{-1}\hat{\mathbf{A}},$ (14)
@@
$\hat{\mathbf{A}}_{ij}$ $=\frac{1}{m}\sum_{k=1}^{m}\psi_{i}(x_{k})\mathcal{L}\psi_{j}(x_{k}),$ $\hat{\mathbf{G}}_{ij}$ $=\frac{1}{m}\sum_{k=1}^{m}\psi_{i}(x_{k})\psi_{j}(x_{k}).$ (15)
These matrices are empirical estimators of the following mass, stiffness, and generator matrices
$\mathbf{A}_{ij}$ $=\left\langle\psi_{i},\,\mathcal{L}\psi_{j}\right\rangle_{\mu}$ $\mathbf{G}_{ij}$ $=\left\langle\psi_{i},\,\psi_{j}\right\rangle_{\mu},$ $\mathbf{L}$ $=\mathbf{G}^{-1}\mathbf{A}.$ (16)
The empirical mass matrix $\hat{\mathbf{G}}$ is often ill-conditioned. A standard approach to circumvent this is to perform a whitening transformation based on removing small eigenvalues
$\widehat{\mathbf{G}}$ $=\mathbf{U}\Sigma\mathbf{U}^{\top},$ $\mathbf{R}$ $=\mathbf{U}\Sigma^{-0.5}\in\mathbb{R}^{n\times r},$ $\hat{\mathbf{L}}_{r}$ $=\mathbf{R}^{\top}\hat{\mathbf{A}}\hskip 2.84544pt\mathbf{R},$ (17)
in which $r\leq n$. Here, $\mathbf{R}$ is a transformation matrix mapping the original basis to the reduced basis
$\mathbf{h}(x)=\mathbf{R}^{\top}\psi(x).$ (18)
Dominant eigenvalues of the generator can be computed by diagonalizing the matrix $\hat{\mathbf{L}}$ or $\hat{\mathbf{L}}_{r}$.
For arbitrary stochastic dynamics, the computation of $\mathbf{A}$ involves a second-order differentiation as shown in Equation (6). However, if the stochastic dynamics are reversible, only first-order derivatives are required to compute the matrix $\mathbf{A}$, as the generator satisfies the following integration-by-parts formula
$\mathbf{A}_{ij}=\left<\psi_{i},\mathcal{L}\psi_{j}\right>_{\mu}=-\frac{1}{2}\int\nabla\psi_{i}\sigma\sigma^{\top}\nabla\psi^{\top}_{j}\,\mathrm{d}\mu.$ (19)
Importantly, if the basis functions are actually defined in a CG space $\hat{\Omega}$, that is $\psi_{i}(x)=\psi_{i}(\xi(x))$, then by the chain rule the matrix $\mathbf{A}$ can be written as
$\begin{split}\mathbf{A}_{ij}&=-\frac{1}{2}\int\nabla_{x}\psi_{i}\sigma\sigma^{\top}\nabla_{x}\psi^{\top}_{j}\,\mathrm{d}\mu=-\frac{1}{2}\int(\nabla_{z}\psi_{i}\nabla_{x}\xi)\sigma\sigma^{\top}(\nabla_{x}\xi^{\top}\nabla_{z}\psi^{\top}_{j})\,\mathrm{d}\mu\\ &=-\frac{1}{2}\int(\nabla_{z}\psi_{i})(\nabla_{x}\xi\sigma\sigma^{\top}\nabla_{x}\xi^{\top})(\nabla_{z}\psi^{\top}_{j})\,\mathrm{d}\mu.\end{split}$ (20)
@@
$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi\sigma\sigma^{\top}\nabla_{x}\xi^{\top}\in\mathbb{R}^{d\times d}$ (21)
@@
2.5 Random Fourier features
The gEDMD algorithm requires choosing a set of basis functions $\psi(x)$. In this work, we use random Fourier features (RFFs), which are defined as
$\psi(x)=\{\cos(\omega_{1}^{\top}x),\,\sin(\omega_{1}^{\top}x),...,\cos(\omega_{n}^{\top}x),\,\sin(\omega_{n}^{\top}x)\}.$ (22)
The vectors $\omega_{1},\ldots,\omega_{n}$ are random frequency vectors drawn from a spectral distribution $\rho$. RFFs provide a low-rank approximation to a reproducing kernel function 29 , and can therefore generate a powerful basis without the need for manual basis set design. The precise relation between kernel-based gEDMD and random features was presented in Ref. 34. In the following applications, we use the spectral measure associated to a Gaussian squared exponential kernel with bandwidth parameter $\gamma$
$k(x_{i},x_{j})=\exp\left(-\frac{\|x_{i}-x_{j}\|)^{2}}{2\gamma^{2}}\right),$ (23)
or to a periodic Gaussian kernel 35 on periodic domains, such as dihedral coordinates.